ANNE FOATA

Death in the Garden:

The Edenic Images in the Fiction of Andrew Lytle

In warning the reader that the family chronicle he is about to read in A Wake for the Living (1975) is the story of an American family’s search for Eden, Andrew Lytle chose to seal by the exemplary story of his own family the major theme of his fiction, proclaiming one last time the importance of the Edenic drive in the making and destiny of his country.

Lytle sees the Edenic myth embedded in the heart of every American in his yearning for innocence and renewal, a dream made impossible, of course, by man’s very expulsion from Paradise, which was also the exact moment of his fall into History. For, in mythical terms, History is the curse visited on postlapsarian mankind, henceforth prey to the slow degradation of time, to the conflicting strivings of the human heart. Lytle in this respect, like so many of his fellow Southern novelists, is no romantic : if man is capable of the best at times, he is also more often and more easily inclined to follow the less strenuous allurements of pettiness and evil. History, as we all know, is of our own making and why should American history escape the common lot ? In Lytle’s vision, the westward drive which brought the Europeans to America and then pushed them to the other side of the continent merely confronted them with themselves, neither better nor worse, whereas in its more brutal form it sent them ravaging and despoiling all over the virgin lands.

In Andrew Lytle’s fiction,^{1} the story of America’s settlement, from the first West wrenched open by the Spaniards’ greed to the many progressive Wests that span the national territory, is the sad story of a large-scale depredation. In the vast continent that opened itself to their heart’s desire, the settlers found unlimited opportunity for the unleashing of their passions and the exercise of unbridled power; abandoning all sense of human contingency they lost themselves in the inflation of their egos and the frenzied pursuit of material ends.

Anne Foata

This is Lytle’s vision of the European venture in the New World, a vision which, at the time of his writing At the Moon’s Inn in the late thirties, prompted him to write a series of novels “progressive in time” whose common theme was to be the different kinds of exploitation perpetrated on the American soil.^{2} With the Spaniards this meant gold and its ruthless pursuit which made them sleep “at the moon’s inn” on a continent unknown to Christian geography, but later it meant the insatiable hunger for land and the aimless wandering of the pioneers engulfed in the “long night” of their soul, and soon after, their fratricidal encounter on the battlefields of the Revolutionary and Civil Wars. Not all of the novels of the projected series appear to have been written, but there is no doubt that the evil for which Lytle sought a name (cf. the title of his third novel A Name for Evil) is the sin of pride which made the Americans believe in a private, privileged communion with God and eventually take themselves for God. Eritis sicut dii, the old Intruder once told our first parents, that of all temptations the temptation most alluring to men whose Faustian appetites may well lead to a worldwide destruction. Always pursuing the dream, the object, or the purpose that lies just ahead of his desire, entrenched in the firm conviction that to pursue is to obtain and to obtain is to be happy and restored in the wholeness of his dream, the American in Lytle’s fiction is “the man with the strained neck, looking always back to Eden” (ANFE 247) and ending up in California or in some swampy jungle on the other side of the world. “To pursue in such a way,” Lytle warns the reader at the beginning of A Wake for the Living, “is to destroy” (Wake 5), a blunt enough statement which the narrator of A Name for Evil endorses even more indisputably when he concludes his considerations on the whole pioneering venture of his countrymen with this chastening piece of truth : “To yearn for the West is simply to yearn for death” (ANFE 169).

And here we might recall what Lytle himself never ceased reminding his readers, that in most of our mythologies, from the Greeks’ to the American Indians’, the West meant death, the grave, the night sea journey, the terror of the abyss which engulfed and sometimes failed to give back; so that nowadays, if the West in general, that is the whole western world, means power and abundance of material goods, the threat of its eternal symbolical meaning still lurks in the western man’s mind.

Death in the Garden

No wonder then that all the blissful Edenic metaphors which blossom in Lytle’s prose reveal themselves full of snares and dangers for his heroes, places of violence and death where the Enemy lies in wait to undo their impossible dream of happiness and renewal. These images — the lush forests, beautiful gardens, fragrant coves — which are so many spatial metonymies of the West and of America, gradually gain in depth while the Edenic quest makes itself felt more and more powerfully as a quest for wholeness, for the recovery of a lost androgynous unity. Thus Lytle’s heroes’ quest becomes a withdrawal from life and his Edenic images as many matricial representations of a dangerous journey back into the womb of the earth and the blissful latency of Chaos or Undifferentiated Matter. It is as if Lytle were reading his Bible backwards: first disposing of Eve to tarry awhile on the androgynous Adam of Genesis 1:27, then discarding Adam with Eden itself, and leaving only the Serpent, no longer the biblical purveyor of apples and temptations, but the slumbering Dragon eating its tail in the waters of Chaos, which in so many cosmogonies preceded the creation of the world, the all-pervading tehom of Genesis 1:2.

“Unity equals zero,” Poe had written in Eureka one century before Lytle dramatized it in A Name for Evil : this equation, if intuitively felt by many artists and by Lytle himself, was the truth which more and more compellingly forced itself upon his creative imagination with the reading of Jung, Neumann and others, and his growing conviction of an ever-recurring archetypal experience within and without the time of man, making of the discovery of America and of its settlement just another phase in the eternal cyclic revolution of death and renewal.

1. The Garden of Delight

In Lytle’s fiction the West, that is the entire American territory in its gradual discovery and settlement, stands for the overall image of Eden, that Paradise of Pleasure that Christopher Columbus dreamed of having opened for the Spaniards’ delight. Then, within this generic metaphor Lytle proceeds, by synecdochic reductions, to confine the Edenic image to a few privileged spots, the forest, a few clearings and coves, and still further reducing, to a tree, a pond, the green shimmer of Creation. At this point the Edenic images proper disappear, absorbed by a quite different kind of representation.

“Think,” de Soto declares to his future companions, the grandees of Spain who will land with him in Florida,

of a land where the air is forever soft, the rivers wholesome from coursing through earth that is never free of gold; think of a climate where a man is rarely driven to close chambers with the cold, nor ever burns his shins before the fire which will make him old before his time by resolving the natural heat of his body. Think of all this and you will be in Florida, the lords of that land. (ATMI 52)

Brave new world indeed, which the discovery of the Indian settlement of Cutifichiqui confirms in its Edenic delight :

There was no fatter land in heathendom, or one more pleasing to the eye. Wide, open forests free of thickets, hummocks of mulberry and oak lying close about the town, and everywhere a great plenty of game… What a land for ease and riches and pleasure. (ATMI 303)

For the weary Spaniards Cutifichiqui is the materialization of the Land of Plenty, of the Paradise of Leisure, and above all, of the Garden of Love where Indian maidens make “bad spinsters” and love is free of sin. Yet there too lurks the serpent and after a while love no longer appears as innocent as it first seemed. A strong intimation of the country’s latent elemental evil is perceived by the narrator when he chances upon a snake in the act of smothering a hawk, then stealing away without even stirring a leave of the underbrush, whereupon he hears the moan of pain of one of his companions in great danger of being emasculated by an Indian damsel.

Three centuries later for the hardy settlers of The Long Night Alabama holds the same promise as Florida once did for the Spaniards. Savannahborn Judge Wilton, who “decided to try his fortune in the West,” dreams of it as “a land of imagination : the vast green forests, the blue mountains, the pleasant fountains, and somewhere a miraculous fortune” (TLN 79). The reader of the novel knows what happens to that dreamer of the American dream : the evil encounters, the “coarse debaucheries of the frontier,” the cheating of the Indians, the “slow death of the great trees,” the large-scale despoiling, and the ensuing moral degeneration. And yet to be sure “he wasn’t by nature a bad man” (TLN 96), only weak, only a man, sinful without reprieve wherever he peddled his dream of renewal. Like de Soto and his soldiers who “came hunting gold and found only themselves” (Wake 50), Judge Wilton, the McIvors and their neighbors only found themselves, the same as ever or just a little worse for the unlimited opportunities that came their way. Robert Penn Warren has a ballad on this theme, and the Billie Pottses must have been legion on the frontier, for

[t]here is always another country and always another place.

There is always another name and another face.

And the name and the face are you, and you

The name and the face.^{3}

Scattered amidst the virgin lands which the pioneers have come to settle, a few “happy places” further concentrate the Edenic mood. Such is the clearing in Georgia where the McIvors make camp on their way to Texas, with spring breaking all at once, the woods full of bird song, the air soft and mild, and a “wet weather spring” not far away. In harmony with their surroundings, the McIvors find themselves in a “frolicsome mood”; there is some fiddling, dancing and courting in the fragrant night. Yet their fate was already sealed and Cameron McIvor’s murder ordained at the darkest of a night no longer kind and welcoming to the heart of the weary.

In The Long Night, forest and night are closely interrelated in their metaphorization of the American wilderness, as images of a dream that has veered to nightmare. How tender indeed the night and friendly the forest to the McIvors’ happy romping near Opelika, and how tutelar and comforting to young love when Damon Harrison and Ruth Weaver “passed… through the great trees… wandering in the strange delight of new, common sensations… and never believing the long night would come to an end” (TLN 169). Yet the promise of life they hold is illusory; everywhere around lurks “danger, invisible and sinister” (TLN 171), and death awaits its victim. The forest which shelters the avenging son, the night which enfolds him in its great tenderness also protect a monomaniac murderer and his reciprocating “secret death with secret death” (TLN 68).

Some time later forest and night seem to breed the violence of the battle of Shiloh and the Civil War, the dead end where the Edenic dream of so many settlers shattered in the nightmare of a fratricide. Synecdoche of the wilderness the Europeans came to settle, the forest in The Long Night is a place of violence and death where the promise of regeneration inscribed in the heart of men stumbles against a dire reality of their own making. The Civil War, in Lytle’s novel as in the novels of many other Southerners, is the image of that stumbling, the young American nation’s definitive Fall from innocence.

The forest imagery, whether or not associated with the night of men’s ordeals, both metonymic representations of the successive Wests that have extended to the other side of the American continent and metaphors of the entire pioneering venture of the Americans, is sustained thoughout Lytle’s fiction. Leaving aside for the moment its deeper mythic abstraction in The Grove and the garden of A Name for Evil, we see it emerge again in the poetic images of The Velvet Horn (set in 1879-80). Here the “deep forest” has just about vanished from the American landscape, leaving only a few patches here and there that are about to be reclaimed for urban and industrial settlements, a “fool’s paradise” soon to be engulfed in the advent of the New South. But how beautiful and fragrant the “April woods” along the old Wilderness trail in Tennessee with their tall trees, clusters of azalea and jasmine, and how “very out of this world” the wood full of honeysuckle around Seven Springs! But if the forest of the Rim, still dense and teeming with game, allows Beverly, Duncan and Julia Cropleigh who wander in it to share in the prelapsarian innocence and unity of all things created, this also is an illusion and a dangerous one at that, for it is a place of transgression where once more the Fall is being enacted. Parcher’s Cove, Beverly’s tamed deer sanctuary, is also the locale of Julia’s seduction, of Duncan’s and Pete Legrand’s murderous encounter, and of the Cropleigh brothers’ “fratricide and suicide and whatever it is to give a leg” (TVH 213). In view of Parcher’s Cove, Tilford Springs offers its soft spongy floor, its moon-drenched atmosphere to other carnal encounters, whence once again, having eaten of the forbidden fruit, the couple is being expelled by the “flaming darts” of the rising sun.

In all the places where the Edenic connotations are emphasized, Lytle took great care that they appear located out of time and space, as intemporal as the drama that is being re-enacted there, gardens of delight for young lovers locked in each other’s arms in that “one posture” “out of all the combinations possible” (TVH 95) in which they strive to restore wholeness.

2. The images of Chaos

At the farthest end of the metonymic process which characterizes Lytle’s Edenic representations one finds the tree : Tree of Life or Knowledge to be sure in many of its manifestations in the novel, but also “seed tree of the world” (TVH 149), “line tree” which “had stood from the time of man” (TVH 3), “one large female tree” topping the triangular cedar glade in Parcher’s Cove (TVH 138), the “high-branched tree growing upwards from the abyss,” with no visible roots or top, which seems to rise from the dull, gelid, opaque surface of the pond barring the entrance to the Cove (TVH 105). Associated with the stagnant water and the “green glaucous gloom” imagery, the tree, however, leaves the Edenic representations proper to enter the realm of the mythic cosmogonies, where, rising from the all-pervading waters and the heavy gloom of Chaos, it is some form of the Cosmic Tree joining earth and firmament together in the act of Creation.

Such indeed does the tree appear to Jack Cropleigh in The Velvet Horn, who with his almost superhuman clairvoyance has much of the seer or the shaman in him :

So, he thought, must the floor of the world have looked on the first day, its patina immaculate in the instancy of creation, that stilled pulsation of the polished film as the instant of time clicked and time began the myriad prick, sprouting… And now within the stilled monotony of his clairvoyance he saw the high-branched tree growing upwards from the abyss. As it grew it was grown, but nowhere could he make out where the roots fed, nor could he measure where the topmost branches paused in their growth. (TVH 104-105)

Another typical example of the subtle shift in time that operates within the Edenic images is to be seen in the scene of Julia’s seduction inside Parcher’s Cove, as relived by her lover Pete Legrand in the “Passionate Husk” section of the novel (TVH 166-168). Parcher’s Cove, it is remembered, is the very image of the Garden in the novel, always beautifully, romantically described whenever it appears, a deer sanctuary to Beverly, Julia’s and Pete’s fragrant love nest; and so indeed it appears to Pete Legrand, out of this world and back in time, “the world simple as the first day.” The scene he evokes here and again later in the novel for Julia’s son’s benefit (TVH 323) is the Genesis temptation scene with Adam, Eve, and the Serpent under “the large female tree” of Knowledge, but the pervading water imagery of the description at the same time pulls it farther backwards in time to the all-pervading waters of Chaos when, surging out of its depth, a Primordial androgynous creature arose to start the work of Creation

In this seasonless grove of cedar he [Pete] had arisen… fluttering the evergreen reflection into which the flames rolled. The air pitched with the color and motion of the sea. Its chaste scent was as astringent as sea water. The laces entwined in his fingers, falling, undulated like seaweed. He crossed over as indifferent as a fish. … Only her hair, aureate, glistened as it floated into the firelight’s surging. … His senses drowned in the fragrance of the rose. (TVH 167-68; my emphasis)

The shift in imagery testifies to a change, or rather a deepening, in Lytle’s apprehension of the Edenic myth. Instead of an untrammeled desire to enjoy, or to engage in wordly pursuits for a maximum of profit and pleasure, the Edenic drive makes itself felt as a yearning for Plenitude which, being impossible in this world, conduces to syndromes of withdrawal and of nostalgia for the blissful state of the Uncreated.

This makes for a quite different interpretation of the Americans’ pioneering venture. Lytle voiced it in his critical study of Caroline Gordon’s novels, writing (in 1949) that behind “the nostalgic effort to return to Eden or some kind of innocence” lay “that temptation of all the most near to a return to Chaos” (Hero 159). That Lytle’s fictional Edenic images can also be apprehended as images of the Undifferentiated Origins of the world thus explains per se why they are so dangerous, so lethal to the heroes of his novels. To yearn for “the prenatural equilibrium of innocence and wholeness,” Lytle writes in “The Working Novelist and the Mythmaking Process,” is to “refus[e] to engage in the cooperating opposites that make life” (Hero 184). Hence the special quality of evil exuded by some of Lytle’s Edenic spots, no longer evil reckoned on a purely moral scale, but some kind of elemental indifference to man, to life itself. There is an intimation of this essential, ontological brand of evil as early as The long Night, when the protagonist in the forest around Shiloh is made aware of its “sullen vicious strength” turned loose on the men who have come to fight in it. In At the Moon’s Inn, the same foreboding as to Florida’s “rightful meaning” and evil comes to Nuño de Tovar at the close of the Edenic Cutifichiqui episode. Florida indeed was the very body of the world. Not through any agent but through its proper self it worked its evil… Had they perchance stumbled upon Eden, abandoned of God, running its unpruned seasons, ignorant of the generations of men, yet throughout all those generations growing heavier with the bloom that cannot die, the decay that cannot live, for the dry rot and the odour of that fruit which blooms and falls, falls and blooms, at the garden’s pole? (ATMI 324)

This image of Florida held as it were in a state of arrested growth, suspended in some kind of eternal summer that never reaps its fruit, is the striking image of life-denying sterility that Lytle will use again in the following novels. For the narrator of A Name for Evil for instance, life, love, should be in the image of an ideal garden “forever in bloom, forever whithholding its seed” (ANFE 254) : “a lovely illusion” indeed, but also so “utterly barren” and finally destructive as his own garden will prove, tranformed by the snow and frost of a winter day into “one blooming mass” of unreal and lethal beauty.

“What a monstrous summertime,” exclaims still another of Lytle’s characters, Jack Cropleigh in The Velvet Horn, “flowerless, fruitless, seedless” (TVH 260). Jack’s hallucinatory vision from down Joe Cree’s open grave into which he has tumbled is that of a green irridescent sphere, of some kind of primal swamp which holds

a lush balance between growth and decay, live eating itself up as it grew, the rolling shine of the green-golden gloom, where nature is never asleep, nor ever awake. A growth so rich it cannot stop to flower. Everything gone to weed. (TVH 260)

His vision, like Tovar’s premonition in Florida, is that of “the body of the world,”

damp, warm from the primal moisture seeping to the molten center, the heat traveling upwards as the rich slime presses down upon the rocky fires. And the sun nowhere present, only this green light inside of which he lay poised. (TVH 260)

It is a representation of the Original Chaos, of the state of Undifferentiated Matter before Creation, neither asleep nor awake, slumbering away in the eternity of a-created Time. Drowsily floating in its midst, like some great primordial fish, or like the fetus inside its mother’s womb, Jack is given the shamanic privilege of witnessing the very act of Creation, when he sees “the lightning horn piercing earth and firmament together,” which, thus fecundated, will bring forth the Primordial androgynous man.

The Lytlean hero’s Edenic nostalgia, after The Long Night, thus makes itself felt more and more powerfully as a yearning for a blissful state of latency, for the “utter rest and oblivion” of nothingness, which is also a strong drive towards annihilation. De Soto’s “conquest” of Florida, as revealed to his ghost in a post mortem shock of recognition, was the vain pursuit of “one bent on his own destruction” (ATMI 397), and Henry Brent’s labors of regeneration in A Name for Evil but thinly disguise his truly Manichaean hatred of life and all created matter. For the Cropleigh siblings in The Velvet Horn, the Edenic dream is a yearning for the impossible state of Wholeness, for a re-union with the other part of their self in the gratifying slumber of the womb, the unconscious. It is a death wish, similar to that of the mystics or to the death ecstasy of Tristan and Isolde, or if not actually for death, then as Erich Neumann explains in The Origins and History of Consciousness which Lytle acknowledged having read during the composition of the novel, at least a very strong wish to return to the latency of fetal life inside the womb, to the “pleromatic stage” of consciousness through what he terms “uroboric incest.”^{4} And here the reader familiar with The Velvet Horn and with Lytle’s essay on the genesis of the novel, “The Working Novelist and the Mythmaking Process”(in Hero), will have recognized the three major clusters of metaphors around which revolve the Edenic images of the novel : the symbols of incest, of androgyny, and of a dangerous passage to the womb, all of them representations of a state of Wholeness which was also the state of Original Chaos.

The images of incest and androgyny are closely related to the one symbol of the Uroboros, one of the many figures — here the Gnostics’ — of the Primordial Monster, Snake or Dragon, lying about the waters of Chaos, eternally impregnating and regenerating itself. As a symbol of Undifferentiation, the Uroboros is contained in the Original Man, Anthropos or Hermaphroditic Adam, “the entire creature,” Lytle writes, “isolated within himself” (Hero 187), whose innocence implies the “stasis” which denies life or creation. For Lytle, “incest” then is the symbol of man’s eternal striving towards the state of Undifferentiated Wholeness, his efforts to become one again through “every kind of human love which the separation had scattered throughout the world” (Hero 188). Paradigmatic of this effort towards Wholeness appears the love between brother and sister, of “the identity-in-essence in the separate-in-form” (TVH 111), but also the all-too-intense love between parents and children, the sexual drive in general, man’s participation mystique with Nature and her beasts, all the forms of intense, exclusive attraction that are represented in the novel. All of them involve violence and death, because “incest” posits a state of nature dangerous for man and therefore “forbidden by divine and human law” (Hero 184), this same state of nature which centuries of civilization have endeavored to subdue or disguise under the cloak of manners and customs, the whole edifice of institutions and conventions devised to restrain the unruly demands of the blood and to check their disruptive power.

In The Velvet Horn, the incestuous drive is everywhere present, even in the animal world of mating turkeys and jackasses; it is enacted however in the legal sense between Duncan and Julia Cropleigh and epitomized in its central character, their brother Jack, whom Lytle has termed the “spiritual hermaphrodite” (Hero 189). Jack is made aware of his own strong incestuous longing during the climactic night of Joe Cree’s wake when right after his vision of the origins of the world he sees himself as part and parcel of an alchemical conjunction of opposites, the desperate attempt at fusion between Duncan as Fire and Julia as Water. In the scene described on page 260, Jack indeed experiences Duncan and Julia’s conjunction as growing out of his own body, “himself on each side of himself moving into himself,” miraculously restored to his Oneness in the middle of the “damp,” “warm,” “green-golden” “irridescent” sphere in which he sees himself floating, “never asleep nor ever awake,” blissfully returned to the latency of the womb.^{5}

The womb imagery thus subsumes almost all the Edenic images of The Velvet Horn, but it is already to be found in A Name for Evil, where Henry Brent’s fascinated stare into the sinkhole in the middle of his estate testifies to his “extravagant longing for abandonment” to “the dangerous pull” of “the black gaping opening” with its promise of “release, escape from the unbearable, the lull of utter rest and oblivion” (ANFE 277-78).

If the images themselves in their moist warm closeness or truly enchanted beauty seem pleasurable and protective enough to the seeker of release and rest, their access, on the contrary, is fraught with dangers and pitfalls which attest to the lethal character of the whole venture. They can be perceived as yet another version of a descensus ad inferos, the descent into the belly of the monster which is also the womb of the earth, the ordeal of so many mythological heroes in their necessary quest for rebirth. In Lytle’s world, however, this ordeal is the end itself, “the participation in withdrawal” (TVH 110) in the great slumbering Void. Consider for instance the two descriptions of Parcher’s Cove in The Velvet Horn on pp. 135-41 and 212-13. The “cleft in the rock” leading to the “jagged entrance” of a “close warm passage” with its “broken and stony bottom” opening on a “dim grey eye” testify to the pitfalls of the journey, which the overall metaphor of the “incestuous tube” transforms into a passage through the castrating obstacles of a vagina dentata. The return to the womb is indeed a mutilating process, as is shown by the spectacle that meets the eye in both descriptions : animals and men either torn apart, physically, or at least spiritually, destroyed.

One also remembers that the beautiful summer garden of the Brents’ estate in A Name for Evil was the cemetery of Major Brent’s six wives and becomes the place both of Henry’s final disintegration and of his wife’s death; and Henry’s difficult progression through the “barrier of thorn trees” and the “tangle of brush and briar” (ANFE 277) towards the sinkhole at the center of the tangled wood again evokes the dangerous labyrinth of the quester’s return to the womb.

It is interesting to note the emergence of these womb images as early as At the Moon’s Inn, where they appear in the night black labyrinth of the Cuban jungle which leads to the clearing where Tovar seduces Doña Eleonora and then in his illumination on the “rightful meaning” of the conquest and the land they have come to conquer. The clearing, the Edenic locale of a quite bewitching scene of seduction, is compared to an “open space” in the middle of a “black waste,” a “savage place,” and access to it is through a “long funnel,” a “labyrinth” through which Tovar has to crawl, “tearing clothes and flesh” (ATMI 89-91), overwhelmed by terror. When Doña Eleonora exclaims “we might be together at the middle of the world,” she merely uses another of the many metaphors of the axis mundi, the geometrical center of the Uncreated Void out of which, in the illud tempus of myth and mythology, soared the creative power of the Cosmic Tree (Mircea Eliade mentions this center of the World as the “creational space” par excellence, 318)

In Tovar’s shock of recognition at the dawn of his “Indian fashion” wedding night, the “garden pole” location of Florida is still another version of the axis mundi in the middle of Chaos, “running its unpruned seasons, ignorant of the generations of men” (ATMI 324). The matricial images which expressed his progress in the Cuban jungle again brush on the surface of his consciousness, with his sensation of being “shut in a narrow room … existing nowhere in space,” and his awareness that Florida was “but a succession of such rooms, each with its corridor, indefinitely multiplied,” the very “body of the world,” the Matrix of the Universe. Tovar’s erotic exertions on his wedding night leave him spent, “consumed” like “dry wood,” and his strenuous journey to the “confines of Christendom” forever stranded in the dark womb of the earth.

The ambiguous character of all the Edenic images in Lytle’s fiction, the ambivalent feelings they arouse in the characters, both a yearning for paradisal bliss and terror of the abyss, their deadly thrust for all those who seek their impossible shelter, in other words, the fascination and terror of the womb, help explain the role and nature of the Lytlean woman. Alluring and destructive, virgin and bride, the mother who lulls and comforts on her life-giving bosom and the death-dealing Gorgon, Eve in Lytle’s Paradise has a precarious stand, wooed and feared, worshipped and derided if not downright denied and destroyed. Eve’s creation, by shattering Adam’s blissful wholeness and making him — unwillingly — participate in the work of Creation, is an obstacle to man’s Edenic nostalgia; and here Jack Cropleigh, the “spiritual hermaphrodite” of The Velvet Horn, very certainly expresses the general feeling of all the male heroes of Lytle’s fiction when he ruefully reflects that Adam must indeed have been tricked by sleep to allow himself to be robbed of his precious rib. The Lytlean man’s inborn terror is undoubtedly to be sucked back “booted and spurred” into “the primordial dark and slime, where the weather’s neither too hot nor too cold, too wet nor dry, just one long growing season” (TVH 90), the image again of the Primordial Womb. His terror, in the last resort, is quite certainly to be denied his place in the universe, to be negated by the uroboric self-regenerating process of Chaos, or swallowed by the great Chthonian Mother in order to be brought forth again in one of those virgin births men do not necessarily appreciate. Jack Cropleigh here is the staunch spokesman of his sex, whose imagination is both fascinated and repelled by the spectacle of the monstrous Ada Rutter cradling her idiot son Othel, paradigmatic of all the incestuous mother-son/lover couples of mythology.

Similarly, to refer to the title of Lytle’s novel, Florida, the whole new world, truly is the “moon’s inn,” the dangerous abode of the Great Goddess of the matriarchal theogonies, who reckoned her tribute “in sweat, hunger and in blood” (ATMI 397), and Diana-like, punished with death all male violators, as Lytle explains in a letter.^{6} Death is the penalty Florida exacts from the Spaniards whose “conquest” appears as the aggressive act of a “purely masculine society” (Hero 158), disguising its destructive will under the cloak of Edenic yearning.Their urge was to conquer, that is to violate, to impose their will on the bountiful virgin nature they might have fertilized instead with their superior skills in the respect and preservation of life, thus committing themselves in the long run to a prolonged dance of self-destruction. Lytle has a parable about this, a story of a marvelous garden and a beautiful woman, and of the destruction of both by a man’s self-deluding dreams of Edenic fulfilment. The story is called A Name for Evil.

3. The Grove, Henry Brent’s fragile crystal Paradise

In A Name for Evil, the Garden metaphor clearly shows through the telling name of the Brent estate, The Grove; it is then further compressed by synecdoche into the flower garden and the wood in the middle of the property which once held the perfection of a Revolutionary major’s pioneering ventures. Henry Brent’s purpose in buying the house and the land which once belonged to his ancestor is to perform a labor of love : the restoration of the dilapidated estate, or rather as he prefers to say, its “regeneration.” It soon becomes evident that by this word he is also thinking of himself, prompted by the Edenic dream of so many of his forebears who had “removed” to the “current West” of their time to start anew a life that had failed elsewhere. More specifically his dream of regeneration concerns his manhood, his abilitly to father a heir after many years of barren wedlock and to give his beloved wife Ellen the status demanded by the American bourgeois standard of living. All this, however, proves to be a self-delusion; Henry’s secret longing is for “release,” “escape,” “the lull of utter rest and oblivion” (ANFE 278), for a return to the state of minimal life which characterizes the state of “uroboric incest,” very akin, we have seen, to the death wish.^{7} He has, in fact, the Manichaeans’ hatred for all created life and his erratic attempts at restoration throughout the novel hardly disguise his one real purpose which is to suppress all life around him, his wife’s, and their longed-for son’s, as once his ancestor had done some three-quarters of a century earlier. The projection of his life-denying impulses on the reappearing specter of the major accounts for the ghost story the novel (also) purports to be, and is quite successfully, despite the parabolic or metaphoric reading Lytle proposes to his reader. I cannot agree, indeed, with Todorov’s contention that any such reading kills the fantastic in a novel of this kind, provided of course it does not prove intrusive.^{8} In A Name for Evil, it does not and the ghosts (for there is a second one) certainly seem real enough (if one may say so), but is is none the less true that the image retained of the American Garden from a certain reading of the novel is an image of death, because those who were sent to “tend” it, as Adam once in Eden, destroyed every semblance of life in their unconscious drive for annihilation.

Henry Brent’s secret loathing of life, which Lytle here posits as inherent to the Americans’ yearning for innocence and Edenic bliss, reveals itself in his conception of ideal love, of the ideal woman and the ideal garden; it materializes in his own garden made ideal and lethal by the frost. In the image of an ideal garden, which, as we have seen, should remain forever in bloom and forever withholding its seed, the woman should be forever the bride on the morning of her nuptials, an “image of inviolability to change” “for all men to behold” (ANFE 184), the impossible fusion of “innocence and desire” (ANFE 270), the recipient of a “grace impossible therefore miraculous, of incorruptible innocence and voluptuous play inexhaustible” (ANFE 295), “pure, trim, and immaculate” (ANFE 184), in short, “an angel” (ANFE 169). A phantasm of man’s immature love, her beauty, “which men of all time have been unable to resist,” calls for his very act of possession, which is an act of will and not of love, an act of darkness, not meant to bear any fruit.

Henry’s recoil (in chapter 26) from Ellen’s kneading the dinner biscuits, “a ritual as old as the world,” which reminds him of the bringing back to life of a drowned body, is revealing, as are his celebration of the tomatoes, the Ancients’ love apples that could also kill, and his calling Ellen Helen. His own vision of himself is that of the perennial bridegroom poised on the narrow edge between the feasting of the eye and the ravishing of the body, the romantic image of eternal, irresponsible youth. To live, to love, to (pro)create, to regenerate, is to partake in “the common filth of living” (ANFE 184), which he of course grandiloquently refuses, and one can safely say that Ellen’s death warrant is signed when she tells him of her unexpected pregnancy, the life rising inside her womb appearing as loathsome to him as the biscuits rising in the oven. The sentence will be executed shortly afterwards in the couple’s garden which winter has made as unreal, as ideal in its fragile crystal beauty as the image that has kept haunting him. Arrested in time, as if suspended in a silent void and transformed into “one blooming mass” of frost, the garden appears frozen into that “unimaginable zero summer” which was T.S. Eliot’s vision of a winter garden in “Little Gidding,” “neither budding nor fading. Not in the scheme of generation.”^{9} This had already been, we remember, Tovar’s vision of Florida, “running its unpruned seasons, oblivious of the generations of men,” before becoming the Cropleighs’ marvelous “seasonless grove” of their own disenchantment. Image of atemporal beauty and sterility, maintained in that one artificial season to which also Lytle refers as “that season which is life-in-death” with its arresting description of “great bloody flowers already drying” and the “blight” everywhere present underneath its breathtaking magnificence (ANFE 276), the Garden, in Lytle’s fiction, is a place that kills.

Through Henry Brent’s destructive self-delusions Lytle allows us to apprehend the nature of the Europeans’ and then the Americans” “westward movement” on the American continent. Henry’s overt, stated purpose in buying his ancestor’s estate is to take stock, to live within a fixed inheritance, or as he says, “to abide” (ANFE 266); he takes pains to make his meaning explicit, adding a few lines further : “I came to a place with a western view” (my emphasis). At this stage of his demented solipsism Henry has quite forgotten what an earlier flash of lucidity had made him perceive, when, reflecting on the many progressive Wests his country has known, he reached the stark conclusion that “[t]o yearn for the West is simply to yearn for death” (ANFE 169).

The “western view” is thus replaced in its “rightful meaning,” disencumbered from the delusions of Edenic renewal, of life-building fantasies. Furthermore, Lytle has Henry Brent re-enact the Major’s selfsame venture, who did remove “to the current West of the time,” and then, in an act of demonic hubris, destroyed his property at the height of its beauty and yield, deeding the house to a spinster daughter and dismissing his unmarried sons with journeymen’s pays.

The garden itself, we remember, was of his making, perfect in its symmetry and all-season flower growth, a manifestation of “utter daring… brilliant imagination… satanic pride” (ANFE 228). It was also the private burying ground of his six wives and mistresses, their tombs laid out in a circle around a springhouse topped by a serpent’s head carved in stone. A place of death with, it its middle, the fountain of rejuvenating water that once so appealed to the Europeans’ imagination (similar to the sinkhole in the middle of the abandoned field that once held the Major’s richest crop), the garden will destroy again, with the Major dressed as a bridegroom (Henri’s major phantasm of innocent love) claiming his ultimate victim in an act of brutal possession which is the grasp of death.

In giving to Henry an ancestor whose very character he embodies and whose very deeds he re-enacts, Lytle supplies the reader with a clue to the true significance of the Americans’ pioneering venture. Prompted by “the hidden nostalgic effort to return to Eden or some state of innocence” and start a new life, it was also a temptation “of all the most near to a return to chaos” (Hero 158), to the blissful inertia of death in the earth’s welcoming womb. This of course makes for a rather unusual unmasking of the secret springs of the Americans’ celebrated love of action and industry, of their practical-minded philosophy, their political endeavors on behalf of progress and democracy. Back in the thirties, in his essay contributed to the Agrarian symposium, Lytle had already warned his countrymen that their senseless pursuit of happiness, the old Edenic dream of mankind made true American-style, was but a sterile and potentially destructive “running around,” which lacked even the logic of a dog chasing its tail (“Tit” 202), a derisory image of the great Uroboros which shapes their secret longing of peaceful slumber and rest.

So in the end one might wonder if the overall metaphor of Lytle’s fiction, its central Edenic image, were not the one contained in the title of his first novel, the “long night” of the soul which is also the timeless night of Chaos, the point of no return of the Americans’ night sea journey, the stumbling block of their dreams of Edenic bliss. And if the new world was to be the old world’s destruction, as Lytle once told his publisher, what a blow finally dealt to the accepted American notion of the frontier to show death embedded in their beautiful virgin Garden and the death wish at the very root of America’s discovery and settlement!

(^{∗}) This paper is an updated version of the article “American History and the Edenic Myth: Andrew Lytle’s Images of the Garden of the World” published in Mythes, Croyances et Religions dans le Monde Anglo-Saxon. Université d’Avignon (Section d’Anglais) 5 (1987): 213–242.

Notes

1The page references of the quotations will be given parenthetically in the text of the article and are those of the following editions of Lytle’s work: The Long Night (1936), abbreviated as TLN, from the 1988 reprint; At the Moon’s Inn, or ATMI, from the original 1941 edition; A Name for Evil, or ANFE, from the 1958 collection A Novel, A Novella and Four Stories; The Velvet Horn, or TVH, from the original 1957 edition; A Wake for the Living, abbreviated as Wake, from the 1975 original. Lytle’s essays which appear in The Hero with the Private Parts will be quoted as Hero with their page reference from the original 1966 edition. His contribution to the Agrarian symposium will appear as (“Tit”) from the 1962 Harper Torchbook edition.

2In a letter to D. L. Chambers, Lytle’s editor at Bobbs-Merrill’s, included in Noel Polk’s bibliography of Lytle’s work in The Mississippi Quarterly 23 (Fall 1970): 453–54.

3“The Ballad of Billie Potts” in Selected Poems, New & Old (New York: Random House, 1966) 230.

4Neumann’s book appeared in Zürich in 1949 and was translated into English in 1954.

5For a more detailed analysis of the symbolism of The Velvet Horn, see my article: “The symbolic structure of Andrew Lytle’s The Velvet Horn,” in The Kentucky Review 5 (Autumn 1983): 16–31.

6Without date or addressee, included in Noel Polk’s bibliography of Lytle’s work, The Mississippi Quarterly 23 (Fall 1970): 455-56.

7Compare Henry Brent’s longing to that of the poet’s soul in D.H. Lawrence’s “The Ship of Death” quoted by Neumann as an illustration of the stage of “uroboric incest” (in The Origins, vol. II, p. 279): “Drift on, drift on, my soul, towards the most pure/most dark oblivion… /the core/of sheer oblivion and utter peace/the womb of silence in the living night/… peace, complete peace!”

8In The Fantastic: A Structural Approach to a Literary Genre (Cleveland: Case Western Reserve University Press, 1973 and Cornell University Press, 1975). Todorov also declares, and I agree, that any second reading, by force less “naive” than the first, is a kind of metareading that jeopardizes if not downright kills the fantastic.

9In Four Quartets (London: Faber & Faber, 1944): 35. Compare Lytle’s description of the garden in chapter 31 with Eliot’s: “Now the hedgerow Is blanched for an hour with transitory blossom Of snow, a bloom more sudden Than that of summer, neither budding nor fading Not in the scheme of generation.”

Works cited

Eliade, Mircea. Traité d’Histoire des Religions. Paris: Payot, 1964; quoted

from the 1974 paperback reprint (Payothèque). Jung, Carl G. Psychology and Alchemy. Vol. 12 of The Collected Works

of Carl G. Jung. Transl. by R.F.C Hull (Bollingen Series XX). Princeton University Press, 1953; revised edition, 1968.

. Mysterium Coniunctionis, Vol. 14 of The Collected Works of C.G. Jung. Transl. by R.F.C. Hull (Bollingen Series XX). Princeton University Press, 1963; 2nd edition 1970.

Lytle, Andrew Nelson. The Long Night. Indianapolis & New York: The Bobbs-Merrill Co., 1936. Rpt. Tuscaloosa: University of Alabama Press, 1988.

. At the Moon’s Inn. Indianapolis & New York: The Bobbs-Merrill Co., 1941. Rpt. Tuscaloosa: University of Alabama Press, 1990.

. A Name for Evil. Indianapolis & New York: The Bobbs-Merrill Co., 1947. Rpt. in A Novel, A Novella and Four Stories. New York: Mc Dowell, Obolensky, 1958.

. The Velvet Horn. New York: Mc Dowell, Obolensky, 1957. Facsimile rpt. Sewanee, Tenn.: The University of the South, 1983.

. A Wake for the Living. New York: Crown, 1975. Rpt. Nashville: J.S. Sanders, 1992.

. The Hero with the Private Parts. Baton Rouge: Louisiana State UP, 1966. Thirteen of its essays are reprinted in Souherners and Europeans. Essays in a Time of Disorder. Baton Rouge: Louisiana State UP, 1988.

. “The Hind Tit.” I’ll Take My Stand. The South and the Agrarian Tradition, by Twelve Southerners. New York: Warner & Bros, 1930. Rpt. New York: Harper & Row (Harper Torchbooks), 1962; revised ed. 1970. Rpt.: Baton Rouge: Louisiana State UP, 1977. Rpt. in From Eden to Babylon. The Social and Political Essays of Andrew Lytle. Washingten, D.C.: Regnery Gateway, 1989.

Neumann, Erich. The Origins and History of Consciousness. Transl. by R.F.C. Hull (Bollingen Series XLII). Princeton University Press, 1954.

. The Great Mother. Transl. by R. Manheim (Bollingen Series XLVII). Princeton University Press, 1955.